OpenStudy (anonymous):

a bag contains 3 blue chips and 2 red chips. Chi and Dolly play the following game. Chi selects one chip at random from the bag. If Chi selects a blue chip, Dolly gives Chi $5. If Chi selects a red chip, Chi gives Dolly $8. Determine Dolly's expectation

OpenStudy (dumbcow):

p(Blue) = 3/5 p(Red) = 2/5 E = (3/5)*-5 + (2/5)*8 E = -3 + 16/5 E = 1/5